### v1

# Momentum Attention: The Physics of In-Context Learning and Spectral Forensics for Mechanistic Interpretability

Abstract
The Mechanistic Interpretability (MI) program has rigorously mapped the Transformer as a precise computational graph (Elhage et al., 2021; Olsson et al., 2022). Building on this solid foundation, we explore the potential of extending this graph with a conservation law and time-varying AC dynamics, thus viewing it as a physical circuit. We introduce Momentum Attention, a symplectic augmentation designed to embed additional physical priors via the kinematic difference operator $p_{t}=q_{t}-q_{t-1}$. Specifically, we implement the symplectic shear transformation $\hat{q}_{t}=q_{t}+\gamma p_{t}$ on queries and keys (leaving values invariant). We identify a fundamental duality between the Symplectic Shear (physics) and the High-Pass Filter (signal processing). This duality is the cornerstone of our contribution: by injecting a kinematic momentum component, we effectively sidestep the topological depth constraint ($L\geq 2$) for induction head formation—a landmark discovery of the MI program for standard Transformers (Olsson et al., 2022). While standard architectures require two layers to derive induction from static positions, our architectural extension grants the model direct access to velocity, enabling Single-Layer Induction while simultaneously unlocking Spectral Forensics via Bode Plots. Addressing the interaction between Low-Pass RoPE (Su and others, 2024) and High-Pass Momentum, we formalize an Orthogonality Theorem, proving the existence of an “Escape Route” where DC (semantic) and AC (mechanistic) signals naturally segregate into orthogonal frequency bands. Validated through 5,100+ controlled experiments, with negative controls (fully documented as an epistemic chronology of discovery in Supplementary Appendices A–R and 27 accompanying Jupyter notebooks with embedded results for reproducibility), our 125M Momentum model exceeds expectations on induction-heavy tasks, while on general-purpose tasks it tracks a 350M baseline within $\sim$2.9% validation loss. We further validate Single-Layer Induction through dedicated associative recall experiments (Addendum to Appendix D), discovering an attenuated scaling law $\gamma^{*}=4.17\times N^{-0.74}$ that establishes momentum-depth fungibility. We humbly offer this framework as a complementary, analytical toolkit for interpretability studies of transformer circuits, connecting Generative AI, Hamiltonian Physics, and Signal Processing.
1 Introduction
The Mechanistic Interpretability (MI) program represents a landmark achievement in deep learning, successfully reverse-engineering the Transformer into a composable computational graph dominated by specific circuits, such as the Induction Heads (Elhage et al., 2021; Olsson et al., 2022; Musaf and others, 2025; Olah et al., 2020; Conmy et al., 2023; Cammarata et al., 2020). These discoveries have provided the community with an invaluable “software” map of In-Context Learning.
Building on top of this success, we propose a complementary extension: imparting time-varying dynamics to these static computational graph circuits. We observe that the standard Transformer architecture displays characteristics of a “DC-Coupled” system. Its attention mechanism operates on positional embeddings that function as static coordinates (Vaswani and others, 2017; Su and others, 2024; Kazemnejad and others, 2024; Press et al., 2021; Shaw et al., 2018). Consequently, the model must often dedicate parameter capacity and topological depth to emulate dynamics that are not explicitly encoded. This is particularly visible in the Induction Head circuit, which typically requires a two-layer composition ($L\geq 2$) to derive the kinematic information necessary for copying and retrieval (see Figure 1 and Appendix B).
We note that the $L\geq 2$ constraint identified by Olsson et al. (2022) is a rigorous and correct consequence of the standard Transformer’s static embedding space. Our work does not contradict this finding; rather, it highlights the architectural trade-off of a “velocity-free” manifold. By introducing Momentum Attention, we utilize a symplectic shear that imparts a physical conservation law alongside time-varying AC dynamics via the Symplectic-High Pass Filter Duality. This architectural extension grants the model direct access to velocity, thereby circumventing the depth constraint and enabling Single-Layer Induction—a capability we validate empirically through dedicated associative recall experiments across multiple network depths (see Figure 5 and Addendum to Appendix D).
This intervention allows us to transform the Transformer from a computational graph to a Physical Circuit. This formulation naturally integrates with the signal processing engineer’s toolkit (Oppenheim and Willsky, 1996; Proakis and Manolakis, 2001), offering Spectral Forensics—the ability to analyze attention heads via Bode plots—as a new lens for analyzing transformer circuits.
Step 1: The Hamiltonian Prior (Symplectic Shear). We define the momentum operator as a backward difference ($p_{t}=q_{t}-q_{t-1}$) and apply a symplectic shear to both the query and key streams. This operation preserves phase space volume (Liouville’s Theorem), satisfying the conservation law required for a stable physical circuit (Li et al., 2018; Chen et al., 2018). We provide the algorithmic implementation in Algorithm 1.
Step 2: The Signal Processing Bridge (Symplectic-Filter Duality). Expanding the momentum term reveals the hidden circuit dynamics. The symplectic shear is mathematically equivalent to a negative feedback loop:
$\hat{q}_{t}=q_{t}+\gamma(q_{t}-q_{t-1})=\underbrace{(1+\gamma)q_{t}}_{\text{Gain}}-\underbrace{\gamma q_{t-1}}_{\text{Feedback}}$ (1)
This derivation reveals a core contribution of our work: the Symplectic-Filter Duality. Equation 1 demonstrates that the physical act of shearing the phase space is mathematically identical to applying a negative feedback loop. This transforms the attention head into a learnable High-Pass Filter, sensitizing it to transitions (AC signals) rather than just states (DC signals) (Oppenheim and Willsky, 1996; Astrom and Murray, 2010).
2 Momentum Attention as a Symplectic Shear
We define the attention mechanism not as a statistical correlation engine, but as a dynamical system evolving in a phase space $\mathcal{M}$ (Goldstein, 2002; Arnol’d, 2013). This section establishes the theoretical guarantees of our method, referencing proofs in Appendices A–C.
2.1 Phase Space Formulation and Uniqueness
Let the input embedding stream be denoted by $X\in\mathbb{R}^{T\times d}$. We define the phase space at time $t$ as the tuple $(q_{t},p_{t})\in\mathcal{M}$.
Definition 2.1 (Kinematic Momentum Operator).
$p_{t}:=\nabla_{t}q_{t}=q_{t}-q_{t-1}$. This operator explicitly captures the local velocity of the semantic trajectory.
Theorem 2.2 (Uniqueness of the Momentum Operator).
The kinematic difference operator $\mathcal{K}(q_{t})=\gamma(q_{t}-q_{t-1})$ is the unique linear operator satisfying: (1) Causality, (2) High-Pass Condition, and (3) Symplectic Consistency.
Proof.
Step 1 (General Form): The most general linear causal operator of history length 1 is $\mathcal{K}(q_{t})=\alpha q_{t}+\beta q_{t-1}$.
Step 2 (High-Pass Constraint): For static input $q_{t}=q_{t-1}=c$, we require $\mathcal{K}(q_{t})=0$. Substituting: $\alpha c+\beta c=(\alpha+\beta)c=0$. Since this must hold for all $c$, we have $\alpha=-\beta$.
Step 3 (Parameterization): Setting $\alpha=\gamma$ yields $\mathcal{K}(q_{t})=\gamma q_{t}-\gamma q_{t-1}=\gamma(q_{t}-q_{t-1})$.
Step 4 (Symplectic Verification): In the $(q,p)$ symplectic basis where $p=q_{t}-q_{t-1}$, the augmentation $q_{\text{new}}=q+\gamma p$ has Jacobian:
$J=\begin{pmatrix}\partial q_{\text{new}}/\partial q&\partial q_{\text{new}}/\partial p\\ \partial p_{\text{new}}/\partial q&\partial p_{\text{new}}/\partial p\end{pmatrix}=\begin{pmatrix}1&\gamma\\ 0&1\end{pmatrix}$ (2)
The determinant $\det(J)=1\cdot 1-\gamma\cdot 0=1$ confirms volume preservation (Liouville’s Theorem). Any non-linear augmentation $q_{\text{new}}=q+f(p)$ would yield $\det(J)=1+\partial f/\partial p\neq 1$ generically, violating symplecticity. ∎∎
2.2 The Symplectic Shear Transformation
The core intervention is the map $\varphi_{\gamma}:\mathcal{M}\to\mathcal{M}$. We explicitly define the transformation matrix $\mathbf{M}$ acting on the phase space vector $(q,p)^{T}$:
$\begin{pmatrix}\hat{q}_{t}\\ \hat{p}_{t}\end{pmatrix}=\mathbf{M}\begin{pmatrix}q_{t}\\ p_{t}\end{pmatrix},\quad\mathbf{M}=\begin{pmatrix}I&\gamma I\\ 0&I\end{pmatrix}$ (3)
Here, $\hat{q}_{t}$ is the momentum-augmented query. The same transformation is applied to the Key stream.
Theorem 2.3 (Preservation of Symplectic Form).
The transformation $\varphi_{\gamma}$ is a symplectic map, preserving the canonical symplectic form $\Omega=\begin{pmatrix}0&I\\ -I&0\end{pmatrix}$.
Proof.
We verify $\mathbf{M}^{T}\Omega\mathbf{M}=\Omega$. Computing:
$\mathbf{M}^{T}\Omega\mathbf{M}=\begin{pmatrix}I&0\\ \gamma I&I\end{pmatrix}\begin{pmatrix}0&I\\ -I&0\end{pmatrix}\begin{pmatrix}I&\gamma I\\ 0&I\end{pmatrix}=\begin{pmatrix}0&I\\ -I&0\end{pmatrix}=\Omega$ (4)
This confirms that $\varphi_{\gamma}$ preserves the symplectic 2-form, ensuring gradient stability during optimization (Goldstein, 2002; Noether, 1918). ∎∎
2.3 The Hamiltonian Shortcut: Single-Layer Induction
Standard transformers require $L\geq 2$ layers for induction because the attention score $A_{t,j}\propto q_{t}^{T}k_{j}$ cannot access $x_{j-1}$. With momentum, this constraint is bypassed.
Theorem 2.4 (Single-Layer Induction Capability).
A single-layer Momentum-Augmented Attention head can implement an approximate Induction Head mechanism without K-composition.
Proof.
The augmented attention score is:
$\text{Score}_{\text{Mom}}=\hat{q}_{t}^{T}k_{j}=((1+\gamma)q_{t}-\gamma q_{t-1})^{T}k_{j}=(1+\gamma)(q_{t}^{T}k_{j})-\gamma(q_{t-1}^{T}k_{j})$ (5)
When momentum is applied symmetrically to both Query and Key streams, the full momentum inner product $\langle p_{t},p_{j}\rangle$ expands as:
$\langle p_{t},p_{j}\rangle=\langle(q_{t}-q_{t-1}),(k_{j}-k_{j-1})\rangle=q_{t}^{T}k_{j}-q_{t}^{T}k_{j-1}-q_{t-1}^{T}k_{j}+q_{t-1}^{T}k_{j-1}$ (6)
The critical term $q_{t-1}^{T}k_{j-1}$ is maximized when the preceding tokens at positions t1t-1 and j1j-1 match. This is precisely the induction condition: if $x_{t}=x_{j}=A$ (current match) AND $x_{t-1}=x_{j-1}$ (previous match), then $q_{t-1}^{T}k_{j-1}\approx\|e_{\text{prev}}\|^{2}$ is large. Thus, a single layer with momentum can match trajectories directly, bypassing K-composition.
This theoretical capability is validated empirically in Figure 5, which presents direct experimental evidence from controlled associative recall experiments (Addendum to Appendix D). In these experiments, a single-layer momentum transformer ($N=1$) achieves 83.4% accuracy on associative recall—a task where the standard transformer ($\gamma=0$) achieves only 1.2% (random chance), confirming that the $L\geq 2$ barrier is genuinely bypassed by the symplectic augmentation rather than merely attenuated.
Algorithm 1 Momentum Augmentation (Symplectic Shear)
0: Embedding stream $X\in\mathbb{R}^{T\times d}$, Coupling $\gamma$
0: Momentum-augmented Queries $Q_{\text{mom}}$, Keys $K_{\text{mom}}$
1: Initialize $q_{0}=X_{0}$, $p_{0}=0$
2:for $t=1$ to $T$ do
3:  $q_{t}\leftarrow X_{t}$
4:  $p_{t}\leftarrow q_{t}-q_{t-1}$ {Kinematic Difference}
5:  $\hat{q}_{t}\leftarrow q_{t}+\gamma p_{t}$ {Symplectic Shear}
6:end for
[…]
9:return $Q_{\text{mom}},K_{\text{mom}}$
2.4 The Expanded Attention Expression
Injecting momentum into both Query ($Q$) and Key ($K$) streams results in an expanded attention score with four distinct interaction terms:
$S=\hat{Q}\hat{K}^{T}=(Q+\gamma\nabla Q)(K+\gamma\nabla K)^{T}=\underbrace{QK^{T}}_{S_{\text{static}}}+\gamma\underbrace{Q(\nabla K)^{T}}_{S_{\text{anticipation}}}+\gamma\underbrace{(\nabla Q)K^{T}}_{S_{\text{drift}}}+\gamma^{2}\underbrace{(\nabla Q)(\nabla K)^{T}}_{S_{\text{kinematic}}}$ (7)
We interpret these four terms as distinct physical circuits: (1) $S_{\text{static}}$ (DC-DC): Standard attention matching static queries to static keys. (2) $S_{\text{anticipation}}$ (DC-AC): Static query, moving key. (3) $S_{\text{drift}}$ (AC-DC): Moving query, static key. (4) $S_{\text{kinematic}}$ (AC-AC): Pure high-pass term matching velocity of query to velocity of key.
2.5 The Transfer Function and Filter Dynamics
To rigorously validate the high-pass nature of the momentum operator, we analyze its transfer function in the Z-domain. The momentum operator applies $y[n]=x[n]+\gamma(x[n]-x[n-1])$. Taking the Z-transform:
$H(z)=(1+\gamma)-\gamma z^{-1}$ (8)
Evaluated on the unit circle $z=e^{j\omega}$:
$H(e^{j\omega})=(1+\gamma)-\gamma e^{-j\omega}$ (9)
At DC ($\omega=0$): $H(1)=1$. At Nyquist ($\omega=\pi$): $H(-1)=1+2\gamma$. Thus, for $\gamma>0$, the magnitude increases with frequency, confirming High-Pass Filter behavior.
Theorem 2.5 (Velocity Transfer Function).
The pure velocity operator $u_{n}=x_{n}-x_{n-1}$ has transfer function $H_{v}(\omega)=1-e^{-j\omega}$ with magnitude $|H_{v}(\omega)|=2|\sin(\omega/2)|$.
Proof.
Computing the squared magnitude:
$|H_{v}(\omega)|^{2}=(1-e^{-j\omega})(1-e^{j\omega})=1-e^{j\omega}-e^{-j\omega}+1=2-2\cos\omega$ (10)
Using the half-angle identity $1-\cos\omega=2\sin^{2}(\omega/2)$:
$|H_{v}(\omega)|^{2}=4\sin^{2}(\omega/2)\implies|H_{v}(\omega)|=2|\sin(\omega/2)|$ (11)
At DC ($\omega=0$): $|H_{v}(0)|=0$ (complete rejection). At Nyquist ($\omega=\pi$): $|H_{v}(\pi)|=2$ (maximum gain). The combined momentum operator interpolates between unity at DC and $(1+2\gamma)$ at Nyquist. ∎∎
2.6 The Orthogonality Theorem (The Escape Route)
We address the interaction between the high-pass momentum signal and the low-pass RoPE.
Theorem 2.6 (Orthogonality of Semantic and Mechanistic Signals).
Given a multi-frequency RoPE basis $\Theta$ and momentum coupling $\gamma$, let $S_{DC}$ be the semantic component (low-frequency) and $S_{AC}$ be the mechanistic component (high-frequency). For $\gamma>\gamma_{c}$, the attention mechanism segregates these signals into orthogonal bands:
$\mathbb{E}[\langle S_{DC},S_{AC}\rangle]\approx\int_{0}^{\pi}H_{LP}(\omega)H_{HP}(\omega)\,d\omega\to 0$ (12)
Proof Sketch. RoPE applies rotation $R_{\theta}(t)=e^{i\theta t}$ with base frequency $\theta$. Low-$\theta$ RoPE acts as a low-pass filter $H_{LP}(\omega)$ concentrated at DC. Momentum acts as high-pass with $|H_{HP}(\omega)|^{2}=4\sin^{2}(\omega/2)$, yielding complete DC rejection at $\omega=0$ and maximum response at Nyquist.
The cross-correlation integral $\int_{0}^{\pi}H_{LP}(\omega)H_{HP}(\omega)\,d\omega$ vanishes when filter supports are disjoint—the “Spectral Escape Route.” Empirically, $\gamma_{c}\approx 0.225$ marks the transition where high-pass gain exceeds cross-term interference, enabling clean signal segregation and the sharp phase transition from random ($\sim$5%) to near-perfect (>>99%) induction accuracy. See Figure 2 and Appendix E for the complete derivation.
2.7 The Placement Corollary (Post-RoPE Validity)
While momentum is applied to both $Q$ and $K$, the location of this injection relative to the RoPE operator is mathematically constrained by physical principles (see Appendix P).
Corollary 2.7 (The Placement Corollary).
To preserve the manifold geometry, the momentum operator must be applied Post-RoPE. Let $R_{t}$ be the RoPE rotation matrix at time $t$.
Post-RoPE (Correct): $\hat{q}=R_{t}q_{t}+\gamma(R_{t}q_{t}-R_{t-1}q_{t-1})$. This correctly computes the kinematic trajectory in the global embedding manifold.
Pre-RoPE (Incorrect): $\hat{q}_{\text{err}}=R_{t}(q_{t}+\gamma(q_{t}-q_{t-1}))$ forces the past token to be rotated by the current frame $R_{t}$, creating a “Frame Mismatch” that destroys relative positional information.
Proof of Non-Commutativity. The error term is:
$\epsilon=P(R(x))-R(P(x))=(R_{t}q_{t}-R_{t-1}q_{t-1})-R_{t}(q_{t}-q_{t-1})=R_{t}q_{t-1}-R_{t-1}q_{t-1}=(R_{t}-R_{t-1})q_{t-1}$ (13)
Using $R_{t}=e^{i\theta}R_{t-1}$, the error magnitude is $|\epsilon|=|1-e^{-i\theta}|\|q_{t-1}\|=2\sin(\theta/2)\|q_{t-1}\|$. This is isomorphic to the classical Coriolis force $F_{C}=-2m(\Omega\times v)$ in rotating frames. At high RoPE frequencies, this error destroys the high-pass signature, yielding $r=0.12$ correlation with theory (vs. $r=0.94$ for correct placement). See Figure 3 and Algorithm 2 for the Spectral Forensics methodology. ∎
2.7.1 Spectral Complementarity: The Conservation Law Consequence
A deeper insight emerges from the conservation-law structure of the symplectic augmentation. Because the momentum operator is derived from a physically grounded Hamiltonian framework—specifically, a volume-preserving shear that satisfies Liouville’s theorem—the resulting high-pass filter $H_{HP}(\omega)$ and the pre-existing low-pass RoPE filter $H_{LP}(\omega)$ are not merely non-interfering but are spectrally complementary. That is, their combined action reconstructs the full input signal:
$|H_{LP}(\omega)|^{2}+|H_{HP}(\omega)|^{2}\approx|H_{\text{input}}(\omega)|^{2}$ (14)
This complementarity is a direct consequence of the symplectic constraint: the shear transformation preserves phase space volume, which in the frequency domain translates to a partition of spectral energy between the DC (semantic) and AC (mechanistic) channels. Unlike an arbitrary high-pass augmentation—which could destructively interfere with RoPE’s positional encoding, amplify noise in overlapping frequency bands, or violate the energy budget of the attention mechanism—the symplectic shear guarantees that spectral energy is redistributed rather than created or destroyed.
This conservation principle provides the theoretical foundation for the dramatic asymmetry observed in Figure 3. When momentum is correctly applied Post-RoPE, the low-pass (RoPE) and high-pass (momentum) filters operate in their respective complementary bands, yielding the clean high-pass Bode signature ($r=0.94$) and the $52.5\%$ performance gain. The filters partition the spectrum faithfully: RoPE preserves semantic content at low frequencies while momentum captures mechanistic transitions at high frequencies, and their sum recovers the complete input information.
In contrast, Pre-RoPE placement violates this complementarity. The Coriolis error (Equation 13) introduces spurious cross-frequency coupling that “smears” spectral energy across bands, destroying the clean partition. The resulting spectral response (left panel of Figure 3) shows no coherent filter structure ($r=0.12$), and the $4.1\%$ regression confirms that the broken complementarity actively degrades performance below the unaugmented baseline. The conservation-law origin of the symplectic shear thus explains not only why correct placement works, but why incorrect placement causes regression: the former preserves the spectral energy partition while the latter violates it.
Figure 1: The Induction Circuit and Phase Transition. (A) Left: Standard two-layer induction head requires Layer 1 (Shift) to pass positional information to Layer 2 (Match), using purely DC signals. Right: Our single-layer Momentum Attention injects dynamic AC signals ($p_{t}=q_{t}-q_{t-1}$) alongside DC signals, enabling Shift+Match in one layer while unlocking Spectral Forensics. (B) Phase transition from Static Regime to Kinematic Regime at $\gamma_{c}\approx 0.225$. Standard transformers require $L\geq 2$ layers; Momentum Attention enables Single-Layer Induction. See Appendices B, D, E and Addendum to Appendix D.
Figure 2: The Orthogonality Theorem: The “Escape Route.” (A) Standard “DC-Coupled” attention processes only semantic states; our “AC-Coupled” Momentum Attention captures both states (DC) and transitions (AC). The Spectral Escape Route emerges when signals occupy orthogonal frequency bands. (B) Empirical frequency response showing DC/AC orthogonality. The critical coupling $\gamma_{c}$ aligns with induction head emergence. See Appendices E, H.
Figure 3: Spectral Forensics: Bode Plot Autopsy. (Top) Kinematic Frame Theory: momentum must be applied Post-RoPE to avoid “Coriolis Error.” (Bottom Left) Pre-RoPE: Frame mismatch destroys spectral signal ($r=0.12$, $4.1\%$ regression). (Bottom Right) Post-RoPE: Clean high-pass signature ($r=0.94$, $52.5\%$ gain). The asymmetry between these outcomes—gain vs. regression, not merely gain vs. parity—is a direct consequence of the spectral complementarity guaranteed by the symplectic conservation law (Section 2.7.1). See Appendices F, P.
Algorithm 2 Spectral Forensics (Bode Extraction)
0: Attention Head Weights $W_{Q},W_{K}$, Frequencies $\omega\in[0,\pi]$
0: Magnitude Response $M(\omega)$
1:for $t=1$ to $T$ do
2:  Generate probe signal $x_{t}=e^{j\omega t}$
3:end for
4: Compute Attention Score $S(\omega)=\text{Attn}(x,x)$
5: Compute Magnitude $M(\omega)=20\log_{10}|S(\omega)|$
6: Plot $M(\omega)$ vs $\omega$ (Bode Plot)
7: Compare with theoretical $H(e^{j\omega})$
8:return $M(\omega)$
3 Empirical Validation
We validate our theoretical claims through an extensive experimental campaign (the “Epistemic Chronology”) comprising over 5,100 controlled runs, detailed fully in Appendices C through R and the Addenda to Appendices B, D, and E. Our validation strategy follows three complementary approaches: (1) spectral forensics to verify filter properties, (2) task dissociation to confirm the $\nabla$-task vs $\int$-task dichotomy, and (3) stress testing to probe the limits of the momentum advantage.
3.1 Spectral Forensics: Theory Meets Experiment
Figure 3 provides a representative validation of spectral forensics in action—the direct empirical measurement of attention head frequency response via Bode plots. This technique, formalized in Algorithm 2, enables us to “autopsy” trained attention heads and verify whether they exhibit the theoretical high-pass characteristics predicted by our framework.
The critical insight from Figure 3 is the dramatic difference between Pre-RoPE and Post-RoPE momentum placement. When momentum is incorrectly applied before RoPE rotation (left panel), the measured frequency response shows essentially no correlation with the theoretical high-pass filter ($r=0.12$), confirming that the “Coriolis Error” described in the Placement Corollary destroys the kinematic signal. In contrast, correct Post-RoPE placement (right panel) yields near-perfect theory-experiment alignment ($r=0.94$), validating both the filter duality and the placement constraint.
[…]
3.2 Task Dissociation: The High-Pass Signature
To isolate the effect of the Momentum Operator on circuit dynamics, we conducted controlled experiments using a 4M parameter proxy model. Our theory predicts that Momentum Attention should excel at “Derivative Tasks” (detecting changes/patterns) while maintaining parity on “Integral Tasks” (accumulating semantic meaning). We expand this analysis to include Chain-of-Thought (CoT) and Multi-Hop reasoning to demonstrate the limits of the momentum prior (Sanford and Hsu, 2024; Wei and others, 2022).
As shown in Table 1, the 4M Momentum model achieves near-perfect accuracy on Single-Layer Induction (98.7%), a task where the standard transformer fails (12.4%) due to the $L\geq 2$ depth constraint (Elhage et al., 2021; Hooper and others, 2024). The pattern is consistent: derivative tasks ($\nabla$-tasks) that require detecting transitions show massive improvements, while integral tasks ($\int$-tasks) that require accumulating information remain at parity. Notably, we observe gains in CoT tasks (Wei and others, 2022; Kojima and others, 2022), suggesting that “kinematic” information aids in tracking reasoning steps—a hybrid behavior consistent with CoT requiring both pattern detection and semantic accumulation.
Table 1: Task Dissociation: $\nabla$-Task vs $\int$-Task Dichotomy. Momentum Attention excels at derivative tasks while maintaining parity on integral tasks. Results from 4M proxy model. See Appendices G, I, J, K, L, M.
Task Type Metric Standard Momentum $\boldsymbol{\Delta}$
Derivative (AC)
   Single-Layer Induction Acc 12.4% 98.7% +86.3%
   Pattern Matching Acc 45.2% 92.1% +46.9%
   Copy/Paste Loss 0.45 0.12 $73\%$
Hybrid (Reasoning)
   Chain-of-Thought (CoT) Acc 62.1% 68.4% +6.3%
   Multi-Hop Reasoning Acc 51.5% 59.2% +7.7%
Integral (DC)
   Language Modeling PPL 18.2 17.8 $2.1\%$
   Semantic Retrieval Acc 76.5% 76.2% $0.3\%$
3.3 Efficiency at Scale: David vs. Goliath
To assess efficacy at scale, we trained a 125M Momentum model and compared it against a 350M Baseline model. While these scales are microscopic by modern SOTA standards, we selected this regime specifically to isolate mechanistic effects and circuit dynamics without the confounding variables inherent in massive-scale training (Kaplan and others, 2020; Hoffmann and others, 2022; Touvron and others, 2023; Chowdhery and others, 2023).
As shown in Table 2, the 125M Momentum model tracks the 350M Baseline within $\sim$2.9% validation loss while using 64% fewer parameters. This validates the “Do No Harm” principle: physics-informed priors can improve parameter efficiency without compromising general capability. See Figure 4 and Appendix R for complete training curves and analysis.
Table 2: David vs. Goliath: Parameter Efficiency at Scale. 125M Momentum model tracks 350M Baseline within $\sim$2.9% validation loss using 64% fewer parameters. Training: 127 GPU-hours, matched hyperparameters. See Appendix R.
Model Params Val Loss
Baseline (Goliath) 350M 2.14
Momentum (David) 125M 2.20
Difference $64\%$ $2.9\%$
3.4 ICL Stress Test: Probing the Limits
Figure 4 presents the most demanding validation of our framework: stress testing In-Context Learning across increasing chain lengths from $L=10$ to $L=50$. This experiment, comprising 2,880 configurations documented in Appendix N, reveals three key insights.
(A) Signal Decay by Depth: Standard attention exhibits characteristic exponential decay in copying fidelity as chain depth increases, consistent with the theoretical $p^{L}$ signal attenuation. Momentum Attention maintains a “Momentum Advantage Zone” where performance degrades more gracefully, achieving linear rather than exponential decay (1cL1-cL).
(B) Theoretical Signal Retention: The middle panel validates our theoretical prediction: the high-pass filter’s DC rejection prevents the accumulation of “semantic drift” that plagues standard attention at long ranges. The momentum term $p_{t}=q_{t}-q_{t-1}$ acts as a differentiator, preserving relative positional information even as absolute positions become unreliable.
(C) Complexity Scaling: Most strikingly, the momentum advantage increases with task complexity. At $L=10$, both architectures perform comparably; by $L=30$, Momentum Attention achieves $52.5\%$ improvement in repetition loss. This scaling behavior suggests that the kinematic prior becomes increasingly valuable precisely when standard attention struggles most—a desirable property for real-world applications requiring long-range pattern matching.
Figure 4: ICL Stress Test. (A) Signal Decay: Standard (red) vs Momentum (blue) across chain depths ($L=30$). (B) Theoretical Retention: exponential decay ($p^{L}$) vs linear decay (1cL1-cL). (C) Complexity Scaling: $52.5\%$ gain from $L=10$ to $L=30$. See Appendices N, O.
3.5 Single-Layer Induction: The Scaling Law
To provide the most direct and rigorous validation of Single-Layer Induction—the central theoretical prediction of our framework (Theorem 2.4)—we conducted a series of dedicated associative recall experiments using controlled synthetic benchmarks, documented in full in the Addendum to Appendix D.
Figure 5 presents the main results from Experiments 16 and 18. Panel (A) shows the definitive evidence for breaking the $N\geq 2$ barrier: a single-layer ($N=1$) momentum transformer achieves 83.4% accuracy on associative recall at $\gamma=4.0$, compared to only 1.2% for the standard transformer ($\gamma=0$)—a $69.5\times$ improvement. The phase transition at $\gamma\approx 1.0$ is clearly visible, with three distinct regimes: sub-critical ($\gamma<0.3$) where the model behaves as a standard transformer, a transition zone ($0.3<\gamma<1.0$) where induction capabilities emerge rapidly, and a saturation regime ($\gamma>4.0$) that reveals a physical limit imposed by the position-momentum uncertainty relation in the embedding space.
Panel (B) reveals the Attenuated Scaling Law: $\gamma^{*}=4.17\times N^{-0.74}$, discovered across network depths $N\in\{1,2,3,4,5,8\}$ with fit quality $R^{2}=0.947$. This power-law relationship establishes a fundamental connection: momentum coupling and network depth are fungible computational resources for induction. The sub-linear exponent ($\alpha=0.74<1$) implies signal attenuation across layers—each additional layer partially absorbs the momentum signal, requiring less coupling to achieve the same induction capability. This relationship provides practical deployment guidance: for a network of depth $N$, the optimal momentum coupling can be predicted a priori from the scaling law, eliminating costly hyperparameter searches.
The scaling law also reveals an important asymmetry: while depth can partially substitute for momentum (deeper networks need less $\gamma$), momentum cannot be fully replaced by depth alone. The standard transformer ($\gamma=0$) fails at associative recall regardless of depth when constrained to a single layer, whereas even modest momentum coupling ($\gamma\approx 1$) unlocks significant capability. This asymmetry reflects the fundamental difference between the “configuration space” (static embeddings) and “phase space” (position + momentum) formulations: the phase space representation provides strictly more information per layer.
Figure 5: The Validated Physics of Symplectic Attention (Experiments 16 & 18). (A) Single-Layer Induction: Breaking the $N\geq 2$ Barrier. The standard transformer ($\gamma=0$, red dashed) achieves only random chance (1.2%), while the momentum transformer (green) reaches 83.4% peak accuracy at $\gamma=4.0$. The phase transition at $\gamma\approx 1.0$ and saturation regime ($\gamma>4.0$, reflecting position-momentum uncertainty) are clearly visible. (B) The Attenuated Scaling Law: $\gamma^{*}=4.17\times N^{-0.74}$. Sub-linear exponent ($\alpha<1$) implies signal attenuation across layers, validating the theoretical prediction that momentum and depth are fungible computational resources. See Addendum to Appendix D for complete experimental details across 270+ configurations.
4 Related Work
Mechanistic Interpretability. Our work builds directly on the foundational circuit analysis of Elhage et al. (2021) and Olsson et al. (2022), who established the rigorous framework for understanding transformers as computational graphs. We extend the geometric analysis of induction heads recently proposed by Musaf and others (2025) and the associative recall analysis by Sanford and Hsu (2024). Our spectral tools complement the automated circuit discovery methods of Conmy et al. (2023) and Olah et al. (2020), as well as the dictionary learning approaches of Bricken and others (2023) and the polysemanticity analysis of Goh and others (2021). The “circuit” metaphor has proven remarkably productive; our contribution extends this metaphor from static computational graphs to dynamic physical circuits.
Physics-Inspired Machine Learning. The intersection of Hamiltonian mechanics and deep learning has been explored extensively in Hamiltonian Neural Networks (Greydanus et al., 2019; Toth et al., 2020) and nonequilibrium thermodynamics (Sohl-Dickstein et al., 2015). We specifically draw inspiration from the renormalization group mappings by Mehta and Schwab (2014) and the statistical mechanics frameworks of Bahri et al. (2020) and Bondesan and Welling (2019). Lagrangian approaches by Cranmer and others (2020) and constrained optimization by Finzi and others (2020) also offer valuable perspectives on conservation laws in learning. Our work differs in applying these principles to the attention mechanism itself rather than to the overall network dynamics.
Signal Processing in Transformers. The role of positional encodings as filters has been studied in RoPE (Su and others, 2024) and ALiBi (Press et al., 2021), as well as Transformer-XL (Dai and others, 2019). Recent work by Kazemnejad and others (2024) and the “KV-Shifting” hypothesis by Hooper and others (2024) align with our kinematic findings. Our spectral forensics approach formalizes these observations using classical signal processing tools (Oppenheim and Willsky, 1996; Proakis and Manolakis, 2001; Kalman, 1960; Feynman et al., 1963). Efficient attention mechanisms like Reformer (Kitaev et al., 2020) and H2O (Zhang and others, 2024) also implicitly touch upon spectral sparsity. The Bode plot methodology we introduce provides a principled way to analyze any attention mechanism’s frequency response.
5 Conclusion
In this work, we have explored the potential of enriching the Transformer’s computational graph with physical conservation laws. By introducing Momentum Attention, we have shown that a simple symplectic augmentation ($p_{t}=q_{t}-q_{t-1}$) can imbue the model with fundamental conservation laws and spectral properties.
This intervention bridges the gap between Hamiltonian mechanics and signal processing. We have demonstrated that the “Symplectic Shear” is mathematically dual to a “High-Pass Filter,” unlocking powerful new tools for analysis—most notably Spectral Forensics. This framework not only explains why the model works (via the Orthogonality Theorem and Escape Routes) but also how to improve it. We humbly offer this work as an invitation for the community to apply the full arsenal of control theory and signal processing to the challenge of interpretability, extending the powerful “circuit” metaphor into the domain of physical dynamics.
Limitations and Future Work. While our experiments validate the theoretical predictions across 5,100+ controlled runs documented in Appendices A–R (with Appendix Q providing complete experimental configuration matrices), several limitations warrant discussion. First, the scale of our models (4M–350M parameters) remains modest compared to frontier systems; extending this framework to billion-parameter scales remains important future work, though the theoretical foundations are scale-agnostic. Second, the optimal momentum coupling $\gamma$ may vary across tasks and architectures—we provide extensive sweeps but acknowledge that adaptive coupling strategies warrant investigation; the attenuated scaling law $\gamma^{*}=4.17\times N^{-0.74}$ discovered in the Addendum to Appendix D provides initial guidance for this. Third, our spectral forensics methodology assumes access to model internals; applying these techniques to black-box models would require additional probing methods.
The symplectic structure naturally suggests extensions to other modalities (vision, audio, video) where temporal dynamics are more explicitly encoded. For video understanding, the kinematic prior could capture motion directly rather than requiring the model to infer it from static frames. For audio, the high-pass filtering interpretation connects to well-understood signal processing principles for speech recognition.
Reproducibility Statement. All experiments are fully reproducible via the 27 Jupyter notebooks provided in the supplementary material (Appendices A–R, including the Addenda to Appendices B, D, and E), with the complete notebook collection also available in the accompanying CODE-NOTEBOOKS-ARXIV.zip archive. Each notebook contains pre-embedded outputs from 5,169+ total experiments, enabling verification without GPU re-execution. Hardware configurations, hyperparameters, random seeds, and training details are documented in exhaustive detail following the “epistemic chronology” philosophy—preserving even productive failures and hypothesis revisions for complete scientific transparency.