### v2

# Momentum Attention: The Physics of In-Context Learning and Spectral Forensics for Mechanistic Interpretability

Abstract
The Mechanistic Interpretability (MI) program has rigorously mapped the Transformer as a precise computational graph (Elhage et al., 2021; Olsson et al., 2022). Building on this solid foundation, we explore the potential of extending this graph with a conservation law and time-varying AC dynamics, thus viewing it as a physical circuit. We introduce Momentum Attention, a symplectic augmentation designed to embed additional physical priors via the kinematic difference operator $p_{t}=q_{t}-q_{t-1}$. Specifically, we implement the symplectic shear transformation $\hat{q}_{t}=q_{t}+\gamma p_{t}$ on queries and keys (leaving values invariant). We identify a fundamental duality between the Symplectic Shear (physics) and the High-Pass Filter (signal processing). This duality is the cornerstone of our contribution: by injecting a kinematic momentum component, we effectively sidestep the topological depth constraint ($L\geq 2$) for induction head formation—a landmark discovery of the MI program for standard Transformers (Olsson et al., 2022). While standard architectures require two layers to derive induction from static positions, our architectural extension grants the model direct access to velocity, enabling Single-Layer Induction while simultaneously unlocking Spectral Forensics via Bode Plots. Addressing the interaction between Low-Pass RoPE (Su and others, 2024) and High-Pass Momentum, we formalize an Orthogonality Theorem, proving the existence of an “Escape Route” where DC (semantic) and AC (mechanistic) signals naturally segregate into orthogonal frequency bands. Validated through 5,100+ controlled experiments, with negative controls (fully documented as an epistemic chronology of discovery in Supplementary Appendices A–R and 27 accompanying Jupyter notebooks with embedded results for reproducibility), our 125M Momentum model exceeds expectations on induction-heavy tasks, while on general-purpose tasks it tracks a 350M baseline within $\sim$2.9% validation loss. We further validate Single-Layer Induction through dedicated associative recall experiments (Addendum to Appendix D), discovering an attenuated scaling law $\gamma^{*}=4.17\times N^{-0.74}$ that establishes momentum-depth fungibility. We humbly offer this framework as a complementary, analytical toolkit for interpretability studies of transformer circuits, connecting Generative AI, Hamiltonian Physics, and Signal Processing.
1 Introduction
The Mechanistic Interpretability (MI) program represents a landmark achievement in deep learning, successfully reverse-engineering the Transformer into a composable computational graph dominated by specific circuits, such as the Induction Heads (Elhage et al., 2021; Olsson et al., 2022; Musaf and others, 2025; Olah et al., 2020; Conmy et al., 2023; Cammarata et al., 2020). These discoveries have provided the community with an invaluable “software” map of In-Context Learning.
Building on top of this success, we propose a complementary extension: imparting time-varying dynamics to these static computational graph circuits. We observe that the standard Transformer architecture displays characteristics of a “DC-Coupled” system. Its attention mechanism operates on positional embeddings that function as static coordinates (Vaswani and others, 2017; Su and others, 2024; Kazemnejad and others, 2024; Press et al., 2021; Shaw et al., 2018). Consequently, the model must often dedicate parameter capacity and topological depth to emulate dynamics that are not explicitly encoded. This is particularly visible in the Induction Head circuit, which typically requires a two-layer composition ($L\geq 2$) to derive the kinematic information necessary for copying and retrieval (see Figure 1 and Appendix B).
We note that the $L\geq 2$ constraint identified by Olsson et al. (2022) is a rigorous and correct consequence of the standard Transformer’s static embedding space. Our work does not contradict this finding; rather, it highlights the architectural trade-off of a “velocity-free” manifold. By introducing Momentum Attention, we utilize a symplectic shear that imparts a physical conservation law alongside time-varying AC dynamics via the Symplectic-High Pass Filter Duality. This architectural extension grants the model direct access to velocity, thereby circumventing the depth constraint and enabling Single-Layer Induction—a capability we validate empirically through dedicated associative recall experiments across multiple network depths (see Figure 5 and Addendum to Appendix D).
This intervention allows us to transform the Transformer from a computational graph to a Physical Circuit. This formulation naturally integrates with the signal processing engineer’s toolkit (Oppenheim and Willsky, 1996; Proakis and Manolakis, 2001), offering Spectral Forensics—the ability to analyze attention heads via Bode plots—as a new lens for analyzing transformer circuits.
Step 1: The Hamiltonian Prior (Symplectic Shear). We define the momentum operator as a backward difference ($p_{t}=q_{t}-q_{t-1}$) and apply a symplectic shear to both the query and key streams. This operation preserves phase space volume (Liouville’s Theorem), satisfying the conservation law required for a stable physical circuit (Li et al., 2018; Chen et al., 2018). We provide the algorithmic implementation in Algorithm 1.
Step 2: The Signal Processing Bridge (Symplectic-Filter Duality). Expanding the momentum term reveals the hidden circuit dynamics. The symplectic shear is mathematically equivalent to a negative feedback loop:
$\hat{q}_{t}=q_{t}+\gamma(q_{t}-q_{t-1})=\underbrace{(1+\gamma)q_{t}}_{\text{Gain}}-\underbrace{\gamma q_{t-1}}_{\text{Feedback}}$ (1)
This derivation reveals a core contribution of our work: the Symplectic-Filter Duality. Equation 1 demonstrates that the physical act of shearing the phase space is mathematically identical to applying a negative feedback loop. This transforms the attention head into a learnable High-Pass Filter, sensitizing it to transitions (AC signals) rather than just states (DC signals) (Oppenheim and Willsky, 1996; Astrom and Murray, 2010).
2 Momentum Attention as a Symplectic Shear
We define the attention mechanism not as a statistical correlation engine, but as a dynamical system evolving in a phase space $\mathcal{M}$ (Goldstein, 2002; Arnol’d, 2013). This section establishes the theoretical guarantees of our method, referencing proofs in Appendices A–C.
2.1 Phase Space Formulation and Uniqueness
Let the input embedding stream be denoted by $X\in\mathbb{R}^{T\times d}$. We define the phase space at time $t$ as the tuple $(q_{t},p_{t})\in\mathcal{M}$.
Definition 2.1 (Kinematic Momentum Operator).
$p_{t}:=\nabla_{t}q_{t}=q_{t}-q_{t-1}$. This operator explicitly captures the local velocity of the semantic trajectory.
Theorem 2.2 (Uniqueness of the Momentum Operator).
The kinematic difference operator $\mathcal{K}(q_{t})=\gamma(q_{t}-q_{t-1})$ is the unique linear operator satisfying: (1) Causality, (2) High-Pass Condition, and (3) Symplectic Consistency.
Proof.
Step 1 (General Form): The most general linear causal operator of history length 1 is $\mathcal{K}(q_{t})=\alpha q_{t}+\beta q_{t-1}$.
Step 2 (High-Pass Constraint): For static input $q_{t}=q_{t-1}=c$, we require $\mathcal{K}(q_{t})=0$. Substituting: $\alpha c+\beta c=(\alpha+\beta)c=0$. Since this must hold for all $c$, we have $\alpha=-\beta$.
Step 3 (Parameterization): Setting $\alpha=\gamma$ yields $\mathcal{K}(q_{t})=\gamma q_{t}-\gamma q_{t-1}=\gamma(q_{t}-q_{t-1})$.
Step 4 (Symplectic Verification): In the $(q,p)$ symplectic basis where $p=q_{t}-q_{t-1}$, the augmentation $q_{\text{new}}=q+\gamma p$ has Jacobian:
$J=\begin{pmatrix}\partial q_{\text{new}}/\partial q&\partial q_{\text{new}}/\partial p\\ \partial p_{\text{new}}/\partial q&\partial p_{\text{new}}/\partial p\end{pmatrix}=\begin{pmatrix}1&\gamma\\ 0&1\end{pmatrix}$ (2)
The determinant $\det(J)=1\cdot 1-\gamma\cdot 0=1$ confirms volume preservation (Liouville’s Theorem). ∎∎
Important Clarification on Non-Linear Shears. A potential objection arises: the linear shear $\Phi_{\text{linear}}:(q,p)\mapsto(q+\gamma p,p)$ is not the only symplectic transformation. Indeed, any map of the form $q^{\prime}=q+f(p)$ where $f(\cdot)$ is differentiable is technically symplectic ($\det J=1$), since the Jacobian of such a map is:
$J_{f}=\begin{pmatrix}1&\partial f/\partial p\\ 0&1\end{pmatrix},\quad\det(J_{f})=1$ (3)
for any differentiable $f$. Therefore, the original Step 4 above does not exclude non-linear shears on symplecticity grounds alone. Instead, we must invoke a stronger physical constraint—Global Lyapunov Stability—to establish uniqueness of the linear form. We develop this argument rigorously below.
2.2 Lyapunov Stability: Why Non-Linear Shear Fails
Beyond symplecticity, we demonstrate that the linear shear is uniquely selected by requiring Global Lyapunov Stability—a necessary condition for the attention mechanism to function as a convergent reasoning process. As shown above, non-linear shears $q^{\prime}=q+f(p)$ can preserve symplecticity. However, we now prove that they cannot simultaneously preserve the convex energy landscape required for stable inference.
Definition 2.3 (Hamiltonian Formulation of Reasoning).
Let the inference process be modeled as a discrete dynamical system minimizing a potential function $V(q)$ (the error landscape), augmented by a kinetic term $T(p)$ (the reasoning momentum). The total Hamiltonian is:
$H(q,p)=T(p)+V(q)$ (4)
The continuous-time equations of motion are given by Hamilton’s equations:
$\dot{q}=\nabla_{p}H=\nabla_{p}T(p),\quad\dot{p}=-\nabla_{q}H=-\nabla_{q}V(q)$ (5)
For the linear shear used in Momentum Attention, the kinetic energy is quadratic: $T_{\text{lin}}(p)=\frac{1}{2}\gamma\|p\|^{2}$.
Theorem 2.4 (Global Stability of Linear Momentum).
If the potential $V(q)$ is convex (standard assumption for local convergence basins), the equilibrium point $(q^{*},0)$ of the system governed by $H_{\text{lin}}$ is globally asymptotically stable under dissipative dynamics.
Proof.
We employ Lyapunov’s Direct Method with candidate function $L(q,p)=H_{\text{lin}}(q,p)=\frac{1}{2}\gamma\|p\|^{2}+V(q)$.
(1) Positive Definiteness: Since $T(p)\geq 0$ and $V(q)$ is locally convex around the minimum, $L(q,p)>0$ for all states except the equilibrium.
(2) Radial Unboundedness: As $\|p\|\to\infty$ or $\|q\|\to\infty$, $L\to\infty$, guaranteeing global coverage.
(3) Orbital Stability: In the conservative case ($\dot{L}=0$), trajectories are closed orbits on level sets of $H$.
(4) Dissipative Convergence: Adding the friction term defined in our architecture ($p_{t+1}=\beta p_{t}+\ldots$), we obtain $\dot{L}<0$.
The Hessian of the kinetic energy is:
$\nabla^{2}_{p}T_{\text{lin}}=\gamma I$ (6)
This is a constant, positive-definite matrix. The geometry of phase space is Euclidean, ensuring straight-line geodesics in momentum space. The system behaves as a Damped Harmonic Oscillator, which is the optimal convergent system. ∎∎
Theorem 2.5 (Instability of Non-Linear Shear).
For non-linear kinetic terms $T(p)$, the Lyapunov candidate $L=H$ fails to guarantee global stability due to loss of convexity in the momentum coordinate.
Proof.
Consider a general non-linear symplectic shear $q^{\prime}=q+f(p)$. This implies a non-quadratic kinetic energy $T_{\text{nonlin}}(p)$ such that $\nabla T=f(p)$. As a concrete example, consider a quartic perturbation commonly encountered in non-linear optics:
$T_{\text{nonlin}}(p)=\frac{1}{2}\|p\|^{2}-\alpha\|p\|^{4}$ (7)
The Hamiltonian becomes $H_{\text{nonlin}}=\frac{1}{2}\|p\|^{2}-\alpha\|p\|^{4}+V(q)$. Compute the Hessian of the kinetic energy with respect to $p$:
$\mathbf{H}_{T}=\nabla^{2}_{p}\left(\frac{1}{2}p^{T}p-\alpha(p^{T}p)^{2}\right)=I-4\alpha\|p\|^{2}I-8\alpha pp^{T}$ (8)
Observe the eigenvalues of $\mathbf{H}_{T}$. For sufficient momentum $\|p\|>\frac{1}{\sqrt{12\alpha}}$, the Hessian becomes negative definite, creating a “Hill-Top” in the kinetic energy landscape. This induces two catastrophic failure modes:
(1) Energy Runaway: If the model accelerates (large $p$), the negative quartic term dominates, causing the Hamiltonian to decrease indefinitely as $\|p\|\to\infty$. The Lyapunov function is no longer radially unbounded.
(2) Chaotic Scattering: Trajectories entering the region where $\nabla^{2}T$ is indefinite become hyperbolic. Small perturbations in initialization lead to exponential divergence of trajectories (positive Lyapunov exponent).
Thus, while non-linear maps may be volume-preserving (symplectic), they destroy the stability basin of the reasoning process. ∎∎
Remark 2.6 (Refined Uniqueness Theorem).
We therefore refine the Uniqueness Theorem as follows. While mathematical symplecticity admits non-linear shears of the form $q^{\prime}=q+f(p)$ for arbitrary differentiable $f$, Physical Robustness selects the linear shear as the unique solution. The linear symplectic shear $\Phi_{\text{linear}}:(q,p)\mapsto(q+\gamma p,p)$ is the only transformation that simultaneously preserves:
  1. 1.
    Phase Space Volume (Liouville’s Theorem, $\det J=1$), and
  2. 2.
    A Convex Energy Landscape (Global Lyapunov Stability, $\nabla^{2}_{p}T=\gamma I\succ 0$).
This dual requirement—symplecticity plus stability—uniquely selects the quadratic kinetic energy $T(p)=\frac{1}{2}\gamma\|p\|^{2}$ and therefore the linear shear. The system is the Hamiltonian analogue of a Damped Harmonic Oscillator, which is widely recognized as the optimal convergent dynamical system in control theory (Astrom and Murray, 2010).
2.3 The Symplectic Shear Transformation
The core intervention is the map $\varphi_{\gamma}:\mathcal{M}\to\mathcal{M}$. We explicitly define the transformation matrix $\mathbf{M}$ acting on the phase space vector $(q,p)^{T}$:
$\begin{pmatrix}\hat{q}_{t}\\ \hat{p}_{t}\end{pmatrix}=\mathbf{M}\begin{pmatrix}q_{t}\\ p_{t}\end{pmatrix},\quad\mathbf{M}=\begin{pmatrix}I&\gamma I\\ 0&I\end{pmatrix}$ (9)
Here, $\hat{q}_{t}$ is the momentum-augmented query. The same transformation is applied to the Key stream.
Theorem 2.7 (Preservation of Symplectic Form).
The transformation $\varphi_{\gamma}$ is a symplectic map, preserving the canonical symplectic form $\Omega=\begin{pmatrix}0&I\\ -I&0\end{pmatrix}$.
Proof.
We verify $\mathbf{M}^{T}\Omega\mathbf{M}=\Omega$. Computing:
$\mathbf{M}^{T}\Omega\mathbf{M}=\begin{pmatrix}I&0\\ \gamma I&I\end{pmatrix}\begin{pmatrix}0&I\\ -I&0\end{pmatrix}\begin{pmatrix}I&\gamma I\\ 0&I\end{pmatrix}=\begin{pmatrix}0&I\\ -I&0\end{pmatrix}=\Omega$ (10)
This confirms that $\varphi_{\gamma}$ preserves the symplectic 2-form, ensuring gradient stability during optimization (Goldstein, 2002; Noether, 1918). ∎∎
2.4 The Hamiltonian Shortcut: Single-Layer Induction
Standard transformers require $L\geq 2$ layers for induction because the attention score $A_{t,j}\propto q_{t}^{T}k_{j}$ cannot access $x_{j-1}$. With momentum, this constraint is bypassed.
Theorem 2.8 (Single-Layer Induction Capability).
A single-layer Momentum-Augmented Attention head can implement an approximate Induction Head mechanism without K-composition.
Proof.
The augmented attention score is:
$\text{Score}_{\text{Mom}}=\hat{q}_{t}^{T}k_{j}=((1+\gamma)q_{t}-\gamma q_{t-1})^{T}k_{j}=(1+\gamma)(q_{t}^{T}k_{j})-\gamma(q_{t-1}^{T}k_{j})$ (11)
When momentum is applied symmetrically to both Query and Key streams, the full momentum inner product $\langle p_{t},p_{j}\rangle$ expands as:
$\langle p_{t},p_{j}\rangle=\langle(q_{t}-q_{t-1}),(k_{j}-k_{j-1})\rangle=q_{t}^{T}k_{j}-q_{t}^{T}k_{j-1}-q_{t-1}^{T}k_{j}+q_{t-1}^{T}k_{j-1}$ (12)
The critical term $q_{t-1}^{T}k_{j-1}$ is maximized when the preceding tokens at positions t1t-1 and j1j-1 match. This is precisely the induction condition: if $x_{t}=x_{j}=A$ (current match) AND $x_{t-1}=x_{j-1}$ (previous match), then $q_{t-1}^{T}k_{j-1}\approx\|e_{\text{prev}}\|^{2}$ is large. Thus, a single layer with momentum can match trajectories directly, bypassing K-composition.
This theoretical capability is validated empirically in Figure 5, which presents direct experimental evidence from controlled associative recall experiments (Addendum to Appendix D). In these experiments, a single-layer momentum transformer ($N=1$) achieves 83.4% accuracy on associative recall—a task where the standard transformer ($\gamma=0$) achieves only 1.2% (random chance), confirming that the $L\geq 2$ barrier is genuinely bypassed by the symplectic augmentation rather than merely attenuated.
Algorithm 1 Momentum Augmentation (Symplectic Shear)
0: Embedding stream $X\in\mathbb{R}^{T\times d}$, Coupling $\gamma$
0: Momentum-augmented Queries $Q_{\text{mom}}$, Keys $K_{\text{mom}}$
1: Initialize $q_{0}=X_{0}$, $p_{0}=0$
2:for $t=1$ to $T$ do
3:  $q_{t}\leftarrow X_{t}$
4:  $p_{t}\leftarrow q_{t}-q_{t-1}$ {Kinematic Difference}
5:  $\hat{q}_{t}\leftarrow q_{t}+\gamma p_{t}$ {Symplectic Shear}
6:end for
[…]
9:return $Q_{\text{mom}},K_{\text{mom}}$
2.5 The Expanded Attention Expression
Injecting momentum into both Query ($Q$) and Key ($K$) streams results in an expanded attention score with four distinct interaction terms:
$S=\hat{Q}\hat{K}^{T}=(Q+\gamma\nabla Q)(K+\gamma\nabla K)^{T}=\underbrace{QK^{T}}_{S_{\text{static}}}+\gamma\underbrace{Q(\nabla K)^{T}}_{S_{\text{anticipation}}}+\gamma\underbrace{(\nabla Q)K^{T}}_{S_{\text{drift}}}+\gamma^{2}\underbrace{(\nabla Q)(\nabla K)^{T}}_{S_{\text{kinematic}}}$ (13)
We interpret these four terms as distinct physical circuits: (1) $S_{\text{static}}$ (DC-DC): Standard attention matching static queries to static keys. (2) $S_{\text{anticipation}}$ (DC-AC): Static query, moving key. (3) $S_{\text{drift}}$ (AC-DC): Moving query, static key. (4) $S_{\text{kinematic}}$ (AC-AC): Pure high-pass term matching velocity of query to velocity of key.
2.6 The Transfer Function and Filter Dynamics
To rigorously validate the high-pass nature of the momentum operator, we analyze its transfer function in the Z-domain. The momentum operator applies $y[n]=x[n]+\gamma(x[n]-x[n-1])$. Taking the Z-transform:
$H(z)=(1+\gamma)-\gamma z^{-1}$ (14)
Evaluated on the unit circle $z=e^{j\omega}$:
$H(e^{j\omega})=(1+\gamma)-\gamma e^{-j\omega}$ (15)
At DC ($\omega=0$): $H(1)=1$. At Nyquist ($\omega=\pi$): $H(-1)=1+2\gamma$. Thus, for $\gamma>0$, the magnitude increases with frequency, confirming High-Pass Filter behavior.
Theorem 2.9 (Velocity Transfer Function).
The pure velocity operator $u_{n}=x_{n}-x_{n-1}$ has transfer function $H_{v}(\omega)=1-e^{-j\omega}$ with magnitude $|H_{v}(\omega)|=2|\sin(\omega/2)|$.
Proof.
Computing the squared magnitude:
$|H_{v}(\omega)|^{2}=(1-e^{-j\omega})(1-e^{j\omega})=1-e^{j\omega}-e^{-j\omega}+1=2-2\cos\omega$ (16)
Using the half-angle identity $1-\cos\omega=2\sin^{2}(\omega/2)$:
$|H_{v}(\omega)|^{2}=4\sin^{2}(\omega/2)\implies|H_{v}(\omega)|=2|\sin(\omega/2)|$ (17)
At DC ($\omega=0$): $|H_{v}(0)|=0$ (complete rejection). At Nyquist ($\omega=\pi$): $|H_{v}(\pi)|=2$ (maximum gain). The combined momentum operator interpolates between unity at DC and $(1+2\gamma)$ at Nyquist. ∎∎
2.7 Rigorous Justification of Small-Signal Analysis
A potential objection to our spectral forensics methodology is that Bode plots are defined for Linear Time-Invariant (LTI) systems, whereas attention is non-linear (Softmax) and time-varying. We address this via the Fréchet Linearization framework, standard in control theory and electrical engineering for analyzing non-linear components (Astrom and Murray, 2010; Oppenheim and Willsky, 1996).
In classical electrical engineering, the frequency response of inherently non-linear components—such as transistors, operational amplifiers, and diodes—is routinely analyzed via Small-Signal Modeling. The key insight is that while these components are globally non-linear, the propagation of small perturbations around a stable operating point occurs in a locally linear regime. We formalize this approach for the attention mechanism.
Definition 2.10 (Small-Signal Approximation).
We decompose the input query vector $q$ into a static operating point $\bar{q}$ and a small perturbation $\delta q(t)$:
$q(t)=\bar{q}+\epsilon\delta q(t),\quad\epsilon\ll 1$ (18)
We seek the linearized transfer function $\mathcal{T}$ such that $\delta y(t)=\mathcal{T}[\delta q(t)]$.
2.7.1 Linearizing the Softmax Operator
The core non-linearity in the attention mechanism is the Softmax function. Let $s=\text{softmax}(x)$, where $x\in\mathbb{R}^{N}$ are the attention logits. The Jacobian matrix $J=\frac{\partial s}{\partial x}$ is given by:
$J_{ij}=s_{i}(\delta_{ij}-s_{j})$ (19)
where $\delta_{ij}$ is the Kronecker delta.
Let the logits be $x=\frac{1}{\sqrt{d}}QK^{T}$. Under the perturbation $Q\to\bar{Q}+\delta Q$, the perturbation in logits is:
$\delta x=\frac{1}{\sqrt{d}}(\delta Q)K^{T}$ (20)
The perturbation in the attention weights $\delta A$ is:
$\delta A\approx J\cdot\delta x=J\cdot\frac{1}{\sqrt{d}}\delta QK^{T}$ (21)
The output of the attention head is $Y=AV$, so the output perturbation is:
$\delta Y=(\delta A)V=\left(J\cdot\frac{1}{\sqrt{d}}\delta QK^{T}\right)V$ (22)
Theorem 2.11 (Local Linearity of Attention).
For a fixed context window (frozen keys $K$ and values $V$), the mapping from a query perturbation $\delta Q$ to the output perturbation $\delta Y$ is a Linear Operator.
Proof.
The expression derived above is of the form $\delta Y=\mathcal{L}(\delta Q)$, where $\mathcal{L}$ involves only matrix multiplications and the constant Jacobian $J$ evaluated at the operating point $\bar{Q}$. Specifically:
$\mathcal{L}(\delta Q)=\left(J\cdot\frac{1}{\sqrt{d}}\delta QK^{T}\right)V$ (23)
Since (i) the Jacobian $J$ is a constant matrix evaluated at the operating point, (ii) $K$ and $V$ are frozen constants for a fixed context, and (iii) matrix multiplication is a linear operation, the composite map $\delta Q\mapsto\delta Y$ is linear. ∎∎
2.7.2 Validation of the Bode Plot Methodology
Remark 2.12 (Spectral Forensics Protocol).
Theorem 2.11 rigorously justifies the “Spectral Forensics” methodology used throughout this paper:
(1) We inject a sinusoidal perturbation: $\delta Q(t)=A\sin(\omega t)$.
(2) By the Local Linearity Theorem, the output $\delta Y(t)$ must be a sinusoid of the same frequency $\omega$, scaled by gain $G(\omega)$ and shifted by phase $\phi(\omega)$.
(3) The ratio $|\delta Y|/|\delta Q|$ is a mathematically valid estimate of the local spectral gain of the attention head at frequency $\omega$.
While the global attention mechanism is non-linear, the propagation of information for small deviations—which correspond to “nuance” or “induction features” in language—occurs in the linear regime. This is directly analogous to how a transistor amplifier is non-linear globally (exhibiting saturation and cutoff regions) but operates linearly for small audio signals within its active region.
Remark 2.13 (Empirical Validation of the Small-Signal Regime).
The validity of the small-signal approximation is empirically confirmed by the Bode plot results in Figure 3. The bottom-right panel of Figure 3 shows the measured frequency response of a trained attention head with correct Post-RoPE momentum placement. The experimental trace (measured from the non-linear attention head) matches the theoretical high-pass filter curve with correlation $r=0.94$. If the small-signal linearization were invalid—i.e., if the non-linear Softmax dynamics dominated even at the perturbation amplitudes used in our probing protocol—the measured response would exhibit non-linear distortion (harmonics, intermodulation products, amplitude-dependent gain shifts) that would destroy the smooth, monotonically increasing Bode signature. The near-perfect agreement between the measured and theoretical traces constitutes strong empirical evidence that the attention mechanism indeed operates in the small-signal linear regime for the perturbation scales relevant to induction feature detection.
2.8 The Orthogonality Theorem (The Escape Route)
We address the interaction between the high-pass momentum signal and the low-pass RoPE.
Theorem 2.14 (Orthogonality of Semantic and Mechanistic Signals).
Given a multi-frequency RoPE basis $\Theta$ and momentum coupling $\gamma$, let $S_{DC}$ be the semantic component (low-frequency) and $S_{AC}$ be the mechanistic component (high-frequency). For $\gamma>\gamma_{c}$, the attention mechanism segregates these signals into orthogonal bands:
$\mathbb{E}[\langle S_{DC},S_{AC}\rangle]\approx\int_{0}^{\pi}H_{LP}(\omega)H_{HP}(\omega)\,d\omega\to 0$ (24)
Proof Sketch. RoPE applies rotation $R_{\theta}(t)=e^{i\theta t}$ with base frequency $\theta$. Low-$\theta$ RoPE acts as a low-pass filter $H_{LP}(\omega)$ concentrated at DC. Momentum acts as high-pass with $|H_{HP}(\omega)|^{2}=4\sin^{2}(\omega/2)$, yielding complete DC rejection at $\omega=0$ and maximum response at Nyquist.
The cross-correlation integral $\int_{0}^{\pi}H_{LP}(\omega)H_{HP}(\omega)\,d\omega$ vanishes when filter supports are disjoint—the “Spectral Escape Route.” Empirically, $\gamma_{c}\approx 0.225$ marks the transition where high-pass gain exceeds cross-term interference, enabling clean signal segregation and the sharp phase transition from random ($\sim$5%) to near-perfect (>>99%) induction accuracy. See Figure 2 and Appendix E for the complete derivation.
2.9 The Placement Corollary (Post-RoPE Validity)
While momentum is applied to both $Q$ and $K$, the location of this injection relative to the RoPE operator is mathematically constrained by physical principles (see Appendix P).
Corollary 2.15 (The Placement Corollary).
To preserve the manifold geometry, the momentum operator must be applied Post-RoPE. Let $R_{t}$ be the RoPE rotation matrix at time $t$.
Post-RoPE (Correct): $\hat{q}=R_{t}q_{t}+\gamma(R_{t}q_{t}-R_{t-1}q_{t-1})$. This correctly computes the kinematic trajectory in the global embedding manifold.
Pre-RoPE (Incorrect): $\hat{q}_{\text{err}}=R_{t}(q_{t}+\gamma(q_{t}-q_{t-1}))$ forces the past token to be rotated by the current frame $R_{t}$, creating a “Frame Mismatch” that destroys relative positional information.
Proof of Non-Commutativity. The error term is:
$\epsilon=P(R(x))-R(P(x))=(R_{t}q_{t}-R_{t-1}q_{t-1})-R_{t}(q_{t}-q_{t-1})=R_{t}q_{t-1}-R_{t-1}q_{t-1}=(R_{t}-R_{t-1})q_{t-1}$ (25)
Using $R_{t}=e^{i\theta}R_{t-1}$, the error magnitude is $|\epsilon|=|1-e^{-i\theta}|\|q_{t-1}\|=2\sin(\theta/2)\|q_{t-1}\|$. This is isomorphic to the classical Coriolis force $F_{C}=-2m(\Omega\times v)$ in rotating frames. At high RoPE frequencies, this error destroys the high-pass signature, yielding $r=0.12$ correlation with theory (vs. $r=0.94$ for correct placement). See Figure 3 and Algorithm 2 for the Spectral Forensics methodology. ∎
2.9.1 Spectral Complementarity: The Conservation Law Consequence
A deeper insight emerges from the conservation-law structure of the symplectic augmentation. Because the momentum operator is derived from a physically grounded Hamiltonian framework—specifically, a volume-preserving shear that satisfies Liouville’s theorem—the resulting high-pass filter $H_{HP}(\omega)$ and the pre-existing low-pass RoPE filter $H_{LP}(\omega)$ are not merely non-interfering but are spectrally complementary. That is, their combined action reconstructs the full input signal:
$|H_{LP}(\omega)|^{2}+|H_{HP}(\omega)|^{2}\approx|H_{\text{input}}(\omega)|^{2}$ (26)
This complementarity is a direct consequence of the symplectic constraint: the shear transformation preserves phase space volume, which in the frequency domain translates to a partition of spectral energy between the DC (semantic) and AC (mechanistic) channels. Unlike an arbitrary high-pass augmentation—which could destructively interfere with RoPE’s positional encoding, amplify noise in overlapping frequency bands, or violate the energy budget of the attention mechanism—the symplectic shear guarantees that spectral energy is redistributed rather than created or destroyed.
This conservation principle provides the theoretical foundation for the dramatic asymmetry observed in Figure 3. When momentum is correctly applied Post-RoPE, the low-pass (RoPE) and high-pass (momentum) filters operate in their respective complementary bands, yielding the clean high-pass Bode signature ($r=0.94$) and the $52.5\%$ performance gain. The filters partition the spectrum faithfully: RoPE preserves semantic content at low frequencies while momentum captures mechanistic transitions at high frequencies, and their sum recovers the complete input information.
In contrast, Pre-RoPE placement violates this complementarity. The Coriolis error (Equation 25) introduces spurious cross-frequency coupling that “smears” spectral energy across bands, destroying the clean partition. The resulting spectral response (left panel of Figure 3) shows no coherent filter structure ($r=0.12$), and the $4.1\%$ regression confirms that the broken complementarity actively degrades performance below the unaugmented baseline. The conservation-law origin of the symplectic shear thus explains not only why correct placement works, but why incorrect placement causes regression: the former preserves the spectral energy partition while the latter violates it.
2.10 From Liouville to Parseval: The Conservation Law Bridge
A rigorous treatment requires connecting two seemingly distinct conservation principles: Liouville’s Theorem (preservation of phase space volume $dq\wedge dp$) and Parseval’s Theorem (preservation of signal energy $\int|F(\omega)|^{2}d\omega$). We argue that in the context of deep learning optimization, Phase Space Collapse is the mechanism of High-Frequency Signal Loss.
2.10.1 Phenomenology vs. First Principles: The FDAM Case
To sharpen this argument, we contrast our approach with the recent Frequency-Dynamic Attention Modulation (FDAM) work by Chen et al. (2025). FDAM observes empirically that standard self-attention acts as a low-pass filter, blurring high-frequency details (e.g., edges in images). To compensate, FDAM employs an ad-hoc inversion to derive a complementary high-pass filter:
$H_{\text{high}}\approx(I-\text{Attn}(x))x$ (27)
This effectively forces the high-frequency components back into the signal path. While effective for dense prediction in Vision Transformers, this approach is phenomenological—it is a “patch” applied to a leaky system, lacking a governing physical law. The high-pass augmentation is engineered from observed behavior rather than derived from first principles, leaving open the possibility of destructive interference, noise amplification, or violation of the attention mechanism’s implicit energy budget.
Our Momentum Attention framework does not patch the leak; it constructs a system that cannot leak. The fundamental distinction is:
FDAM (Phenomenological): Observe low-pass behavior $\to$ invert to create high-pass $\to$ hope for consistency.
Momentum Attention (First Principles): Impose symplecticity ($\det J=1$) $\to$ conservation law forbids rank collapse $\to$ high-pass filter emerges as a mathematical consequence.
2.10.2 The Bridge: Phase Space Volume $\Leftrightarrow$ Signal Energy
Definition 2.16 (Phase Space Collapse).
Rank Collapse occurs when a Transformer layer projects embeddings into a lower-dimensional subspace, compressing phase space volume. This compression preferentially eliminates high-frequency components because they typically reside in the tail of the singular value spectrum.
By enforcing symplecticity ($\det J=1$), we forbid rank collapse. By maintaining the full volume of the container (Phase Space), we ensure the content (Signal Energy) is preserved. The connection proceeds as follows:
Step 1 (Liouville). The symplectic shear preserves phase space volume: $\int dq\,dp=\int d\hat{q}\,d\hat{p}$. This prevents the embedding manifold from collapsing onto a lower-dimensional subspace.
Step 2 (Rank Preservation). Volume preservation implies that the Jacobian of the transformation has unit determinant at every point, which in turn implies that no singular value can approach zero. The transformation maintains full rank.
Step 3 (Spectral Consequence). Full-rank preservation ensures that all frequency components of the signal—including the high-frequency “tail” that is most vulnerable to rank collapse—are retained through the transformation.
Step 4 (Parseval). Since the signal is preserved at full rank, the total signal energy $\int|F(\omega)|^{2}d\omega$ is conserved. The conservation of phase space volume (Liouville) thus implies conservation of signal energy (Parseval) in the context of attention operations.
2.10.3 Spread Spectrum via Symplectic Structure
Theorem 2.17 (Spread Spectrum via Symplectic Structure).
The momentum operator induces orthogonal signal channels for semantic (DC) and mechanistic (AC) content, satisfying:
$y_{\text{total}}(t)=\underbrace{y_{\text{sem}}(t)}_{\text{Low Freq}}+\gamma\underbrace{z(t)}_{\text{High Freq}}$ (28)
where z(t)=y(t)y(t1)z(t)=y(t)-y(t-1) is the momentum signal with transfer function $H_{\text{mom}}(\omega)=1-e^{-j\omega}$.
Proof.
The momentum operator in the Z-domain (discrete frequency domain) is:
$z_{t}=y_{t}-y_{t-1}\implies Z(z)=Y(z)(1-z^{-1})$ (29)
The frequency response, evaluated at $z=e^{j\omega}$, is $H_{\text{mom}}(\omega)=1-e^{-j\omega}$. The magnitude response is:
$|H_{\text{mom}}(\omega)|^{2}=|1-(\cos\omega-j\sin\omega)|^{2}=(1-\cos\omega)^{2}+\sin^{2}\omega=2-2\cos\omega=4\sin^{2}\left(\frac{\omega}{2}\right)$ (30)
This is a canonical high-pass filter with null at DC ($\omega=0$) and peak at Nyquist ($\omega=\pi$).
The interference between semantic and mechanistic signals is measured by their inner product (via Parseval’s identity):
$\langle y_{\text{sem}},\gamma z\rangle=\frac{1}{2\pi}\int_{-\pi}^{\pi}\hat{Y}_{\text{sem}}(\omega)\cdot\gamma\hat{Z}(\omega)\,d\omega\propto\int_{-\pi}^{\pi}\hat{Y}_{\text{sem}}(\omega)\cdot\sin\left(\frac{\omega}{2}\right)d\omega$ (31)
Support Separation: Semantic evolution in text is slow (long-range dependencies), so $\hat{Y}_{\text{sem}}(\omega)$ has compact support near $\omega\approx 0$.
Filter Rejection: The momentum filter term $\sin(\omega/2)$ vanishes at $\omega=0$.
Integral Vanishing: The product of a function concentrated at 0 and a function that is 0 at 0 yields $|\int\hat{Y}_{\text{sem}}\hat{Z}|\leq\epsilon$.
This proves that mechanistic signals (carried by momentum) travel on a channel orthogonal to semantic signals—the definition of Spread Spectrum technology (CDMA), derived purely from the symplectic ansatz. ∎∎
Remark 2.18 (FDAM vs. Momentum: A Plumbing Analogy).
The distinction between the FDAM approach and our framework can be summarized via a plumbing analogy. FDAM observes that the pipe (attention) is clogged (low-pass filtering removes high-frequency details) and applies a plunger (attention inversion) to unblock it. Our Momentum Attention, by contrast, constructs a pipe that cannot clog in the first place: the symplectic constraint ($\det J=1$) provides a physical law guaranteeing that phase space volume—and hence signal energy across all frequency bands—is preserved throughout the transformation. The practical consequence is that FDAM’s correction may introduce artifacts (noise amplification, destructive interference) in regimes where the phenomenological inversion breaks down, whereas our conservation-law-based approach provides structural guarantees by construction.
Figure 1: The Induction Circuit and Phase Transition. (A) Left: Standard two-layer induction head requires Layer 1 (Shift) to pass positional information to Layer 2 (Match), using purely DC signals. Right: Our single-layer Momentum Attention injects dynamic AC signals ($p_{t}=q_{t}-q_{t-1}$) alongside DC signals, enabling Shift+Match in one layer while unlocking Spectral Forensics. (B) Phase transition from Static Regime to Kinematic Regime at $\gamma_{c}\approx 0.225$. Standard transformers require $L\geq 2$ layers; Momentum Attention enables Single-Layer Induction. See Appendices B, D, E and Addendum to Appendix D.
Figure 2: The Orthogonality Theorem: The “Escape Route.” (A) Standard “DC-Coupled” attention processes only semantic states; our “AC-Coupled” Momentum Attention captures both states (DC) and transitions (AC). The Spectral Escape Route emerges when signals occupy orthogonal frequency bands. (B) Empirical frequency response showing DC/AC orthogonality. The critical coupling $\gamma_{c}$ aligns with induction head emergence. See Appendices E, H.
Figure 3: Spectral Forensics: Bode Plot Autopsy. (Top) Kinematic Frame Theory: momentum must be applied Post-RoPE to avoid “Coriolis Error.” (Bottom Left) Pre-RoPE: Frame mismatch destroys spectral signal ($r=0.12$, $4.1\%$ regression). (Bottom Right) Post-RoPE: Clean high-pass signature ($r=0.94$, $52.5\%$ gain). The asymmetry between these outcomes—gain vs. regression, not merely gain vs. parity—is a direct consequence of the spectral complementarity guaranteed by the symplectic conservation law (Section 2.9.1). See Appendices F, P.
Algorithm 2 Spectral Forensics (Bode Extraction)
0: Attention Head Weights $W_{Q},W_{K}$, Frequencies $\omega\in[0,\pi]$
0: Magnitude Response $M(\omega)$
1:for $t=1$ to $T$ do
2:  Generate probe signal $x_{t}=e^{j\omega t}$
3:end for
4: Compute Attention Score $S(\omega)=\text{Attn}(x,x)$
5: Compute Magnitude $M(\omega)=20\log_{10}|S(\omega)|$
6: Plot $M(\omega)$ vs $\omega$ (Bode Plot)
7: Compare with theoretical $H(e^{j\omega})$
8:return $M(\omega)$
3 Empirical Validation
We validate our theoretical claims through an extensive experimental campaign (the “Epistemic Chronology”) comprising over 5,100 controlled runs, detailed fully in Appendices C through R and the Addenda to Appendices B, D, and E. Our validation strategy follows three complementary approaches: (1) spectral forensics to verify filter properties, (2) task dissociation to confirm the $\nabla$-task vs $\int$-task dichotomy, and (3) stress testing to probe the limits of the momentum advantage.
3.1 Spectral Forensics: Theory Meets Experiment
Figure 3 provides a representative validation of spectral forensics in action—the direct empirical measurement of attention head frequency response via Bode plots. This technique, formalized in Algorithm 2, enables us to “autopsy” trained attention heads and verify whether they exhibit the theoretical high-pass characteristics predicted by our framework.
The critical insight from Figure 3 is the dramatic difference between Pre-RoPE and Post-RoPE momentum placement. When momentum is incorrectly applied before RoPE rotation (left panel), the measured frequency response shows essentially no correlation with the theoretical high-pass filter ($r=0.12$), confirming that the “Coriolis Error” described in the Placement Corollary destroys the kinematic signal. In contrast, correct Post-RoPE placement (right panel) yields near-perfect theory-experiment alignment ($r=0.94$), validating both the filter duality and the placement constraint.
[…]
3.2 Task Dissociation: The High-Pass Signature
To isolate the effect of the Momentum Operator on circuit dynamics, we conducted controlled experiments using a 4M parameter proxy model. Our theory predicts that Momentum Attention should excel at “Derivative Tasks” (detecting changes/patterns) while maintaining parity on “Integral Tasks” (accumulating semantic meaning). We expand this analysis to include Chain-of-Thought (CoT) and Multi-Hop reasoning to demonstrate the limits of the momentum prior (Sanford and Hsu, 2024; Wei and others, 2022).
As shown in Table 1, the 4M Momentum model achieves near-perfect accuracy on Single-Layer Induction (98.7%), a task where the standard transformer fails (12.4%) due to the $L\geq 2$ depth constraint (Elhage et al., 2021; Hooper and others, 2024). The pattern is consistent: derivative tasks ($\nabla$-tasks) that require detecting transitions show massive improvements, while integral tasks ($\int$-tasks) that require accumulating information remain at parity. Notably, we observe gains in CoT tasks (Wei and others, 2022; Kojima and others, 2022), suggesting that “kinematic” information aids in tracking reasoning steps—a hybrid behavior consistent with CoT requiring both pattern detection and semantic accumulation.
Table 1: Task Dissociation: $\nabla$-Task vs $\int$-Task Dichotomy. Momentum Attention excels at derivative tasks while maintaining parity on integral tasks. Results from 4M proxy model. See Appendices G, I, J, K, L, M.
Task Type Metric Standard Momentum $\boldsymbol{\Delta}$
Derivative (AC)
   Single-Layer Induction Acc 12.4% 98.7% +86.3%
   Pattern Matching Acc 45.2% 92.1% +46.9%
   Copy/Paste Loss 0.45 0.12 $73\%$
Hybrid (Reasoning)
   Chain-of-Thought (CoT) Acc 62.1% 68.4% +6.3%
   Multi-Hop Reasoning Acc 51.5% 59.2% +7.7%
Integral (DC)
   Language Modeling PPL 18.2 17.8 $2.1\%$
   Semantic Retrieval Acc 76.5% 76.2% $0.3\%$
3.3 Efficiency at Scale: David vs. Goliath
To assess efficacy at scale, we trained a 125M Momentum model and compared it against a 350M Baseline model. While these scales are microscopic by modern SOTA standards, we selected this regime specifically to isolate mechanistic effects and circuit dynamics without the confounding variables inherent in massive-scale training (Kaplan and others, 2020; Hoffmann and others, 2022; Touvron and others, 2023; Chowdhery and others, 2023).
As shown in Table 2, the 125M Momentum model tracks the 350M Baseline within $\sim$2.9% validation loss while using 64% fewer parameters. This validates the “Do No Harm” principle: physics-informed priors can improve parameter efficiency without compromising general capability. See Figure 4 and Appendix R for complete training curves and analysis.
Table 2: David vs. Goliath: Parameter Efficiency at Scale. 125M Momentum model tracks 350M Baseline within $\sim$2.9% validation loss using 64% fewer parameters. Training: 127 GPU-hours, matched hyperparameters. See Appendix R.
Model Params Val Loss
Baseline (Goliath) 350M 2.14
Momentum (David) 125M 2.20
Difference $64\%$ $2.9\%$
3.4 ICL Stress Test: Probing the Limits
Figure 4 presents the most demanding validation of our framework: stress testing In-Context Learning across increasing chain lengths from $L=10$ to $L=50$. This experiment, comprising 2,880 configurations documented in Appendix N, reveals three key insights.
(A) Signal Decay by Depth: Standard attention exhibits characteristic exponential decay in copying fidelity as chain depth increases, consistent with the theoretical $p^{L}$ signal attenuation. Momentum Attention maintains a “Momentum Advantage Zone” where performance degrades more gracefully, achieving linear rather than exponential decay (1cL1-cL).
(B) Theoretical Signal Retention: The middle panel validates our theoretical prediction: the high-pass filter’s DC rejection prevents the accumulation of “semantic drift” that plagues standard attention at long ranges. The momentum term $p_{t}=q_{t}-q_{t-1}$ acts as a differentiator, preserving relative positional information even as absolute positions become unreliable.
(C) Complexity Scaling: Most strikingly, the momentum advantage increases with task complexity. At $L=10$, both architectures perform comparably; by $L=30$, Momentum Attention achieves $52.5\%$ improvement in repetition loss. This scaling behavior suggests that the kinematic prior becomes increasingly valuable precisely when standard attention struggles most—a desirable property for real-world applications requiring long-range pattern matching.
Figure 4: ICL Stress Test. (A) Signal Decay: Standard (red) vs Momentum (blue) across chain depths ($L=30$). (B) Theoretical Retention: exponential decay ($p^{L}$) vs linear decay (1cL1-cL). (C) Complexity Scaling: $52.5\%$ gain from $L=10$ to $L=30$. See Appendices N, O.
3.5 Single-Layer Induction: The Scaling Law
To provide the most direct and rigorous validation of Single-Layer Induction—the central theoretical prediction of our framework (Theorem 2.8)—we conducted a series of dedicated associative recall experiments using controlled synthetic benchmarks, documented in full in the Addendum to Appendix D.
Figure 5 presents the main results from Experiments 16 and 18. Panel (A) shows the definitive evidence for breaking the $N\geq 2$ barrier: a single-layer ($N=1$) momentum transformer achieves 83.4% accuracy on associative recall at $\gamma=4.0$, compared to only 1.2% for the standard transformer ($\gamma=0$)—a $69.5\times$ improvement. The phase transition at $\gamma\approx 1.0$ is clearly visible, with three distinct regimes: sub-critical ($\gamma<0.3$) where the model behaves as a standard transformer, a transition zone ($0.3<\gamma<1.0$) where induction capabilities emerge rapidly, and a saturation regime ($\gamma>4.0$) that reveals a physical limit imposed by the position-momentum uncertainty relation in the embedding space.
Panel (B) reveals the Attenuated Scaling Law: $\gamma^{*}=4.17\times N^{-0.74}$, discovered across network depths $N\in\{1,2,3,4,5,8\}$ with fit quality $R^{2}=0.947$. This power-law relationship establishes a fundamental connection: momentum coupling and network depth are fungible computational resources for induction. The sub-linear exponent ($\alpha=0.74<1$) implies signal attenuation across layers—each additional layer partially absorbs the momentum signal, requiring less coupling to achieve the same induction capability. This relationship provides practical deployment guidance: for a network of depth $N$, the optimal momentum coupling can be predicted a priori from the scaling law, eliminating costly hyperparameter searches.
The scaling law also reveals an important asymmetry: while depth can partially substitute for momentum (deeper networks need less $\gamma$), momentum cannot be fully replaced by depth alone. The standard transformer ($\gamma=0$) fails at associative recall regardless of depth when constrained to a single layer, whereas even modest momentum coupling ($\gamma\approx 1$) unlocks significant capability. This asymmetry reflects the fundamental difference between the “configuration space” (static embeddings) and “phase space” (position + momentum) formulations: the phase space representation provides strictly more information per layer.
Figure 5: The Validated Physics of Symplectic Attention (Experiments 16 & 18). (A) Single-Layer Induction: Breaking the $N\geq 2$ Barrier. The standard transformer ($\gamma=0$, red dashed) achieves only random chance (1.2%), while the momentum transformer (green) reaches 83.4% peak accuracy at $\gamma=4.0$. The phase transition at $\gamma\approx 1.0$ and saturation regime ($\gamma>4.0$, reflecting position-momentum uncertainty) are clearly visible. (B) The Attenuated Scaling Law: $\gamma^{*}=4.17\times N^{-0.74}$. Sub-linear exponent ($\alpha<1$) implies signal attenuation across layers, validating the theoretical prediction that momentum and depth are fungible computational resources. See Addendum to Appendix D for complete experimental details across 270+ configurations.
4 From Computation Graphs to Physical Circuits: Resolving Dynamic Phenomena in Mechanistic Interpretability
The Mechanistic Interpretability (MI) program has achieved remarkable success in reverse-engineering the Transformer as a precise computational graph (Elhage et al., 2021; Olsson et al., 2022; Olah et al., 2020; Conmy et al., 2023). The “circuit” metaphor—where attention heads and MLPs are identified as discrete, composable modules implementing specific functions—has provided the community with an invaluable roadmap for understanding In-Context Learning.
However, recent empirical findings have identified dynamic phenomena that challenge the static circuit picture. We respectfully suggest that these phenomena are not failures of the MI framework, but rather indications that the computational graph can be productively enriched with physical structure. Specifically, we propose that augmenting the static circuit with a conservation law (Liouville’s Theorem) and time-varying AC dynamics (the momentum operator) transforms it into a physical circuit—a representation that naturally encompasses dynamic effects while preserving all the insights of the original static analysis.
4.1 The Hydra Effect: Self-Repair via Conservation Laws
McGrath et al. (2023) identified the “Hydra Effect,” a striking phenomenon where ablating a specific attention head causes other, previously dormant heads to spontaneously take over its function. This emergent self-repair behavior—named for the mythological creature that regrows severed heads—has profound implications for circuit-level attribution: it suggests that “circuits” are not rigid wires with fixed functions, but fluid functional pathways that can dynamically redistribute computation.
Our phase space framework provides a natural explanation for the Hydra Effect via the conservation laws inherent in the symplectic structure.
Remark 4.1 (Conservation Law Interpretation of the Hydra Effect).
In a symplectic system governed by Liouville’s Theorem, phase space volume is conserved regardless of the specific channel through which the system evolves. By analogy, consider a fluid dynamic system: the total mass flow is preserved regardless of which specific pipe carries the fluid. If one pipe is blocked (ablated), the fluid redistributes through the remaining pipes to satisfy the conservation constraint.
In the attention mechanism, the “High-Pass Filter” function—the ability to detect semantic transitions and perform induction—is a property of the layer’s dynamics, not of any specific neuron or head. This function is determined by the spectral requirements of the task: if the task demands detection of high-frequency transition patterns (AC signals), the system must satisfy this spectral constraint. When one attention head is ablated, the gradient descent process during training (or, in the case of instant self-repair, the residual stream’s natural redistribution dynamics) re-optimizes the $\gamma$ coupling of the remaining heads to collectively satisfy the spectral requirements.
Mathematically, the key insight is that the symplectic constraint $\det J=1$ is a global property of the layer, not a property of individual heads. If one head’s contribution to the symplectic shear is removed, the remaining heads adjust their coupling constants to maintain the aggregate conservation law. This is directly analogous to Kirchhoff’s Current Law in electrical circuits: the total current entering a node must equal the total current leaving, regardless of which specific branch carries how much current. The Hydra Effect is thus the neural network’s version of current redistribution in a physical circuit.
Definition 4.2 (Spectral Budget Conservation).
For a layer with $H$ attention heads, each with coupling $\gamma_{h}$, the aggregate spectral transfer function is:
$H_{\text{layer}}(\omega)=\sum_{h=1}^{H}\alpha_{h}\cdot H_{h}(\omega;\gamma_{h})$ (32)
where $\alpha_{h}$ are the residual stream mixing coefficients. The conservation law implies that ablating head $h^{*}$ induces a redistribution $\gamma_{h}\to\gamma^{\prime}_{h}$ for $h\neq h^{*}$ such that $H_{\text{layer}}(\omega)$ is approximately preserved across the task-relevant frequency band.
4.2 Superposition and Polysemanticity: Frequency-Domain Resolution
Elhage et al. (2022) describe “Superposition,” a phenomenon where individual neurons encode multiple disparate features simultaneously—a condition termed “polysemanticity” that significantly complicates circuit analysis. From the perspective of the static circuit picture, superposition appears as an irreducible interference pattern: multiple features sharing the same neuron seems to violate the principle that circuits should be cleanly decomposable into interpretable components.
We propose that this apparent confusion arises from analyzing the system exclusively in the spatial domain (which neurons activate) while ignoring the frequency domain (at what temporal scale the activations vary). By applying Spectral Forensics, we can resolve superposed features in the frequency domain, revealing that they occupy orthogonal spectral bands:
DC Band (Low Frequency): Carries static semantic content—the “meaning” of the current context (e.g., “The cat is on the…”). This information varies slowly across the token sequence and occupies the low-frequency spectral band.
AC Band (High Frequency): Carries mechanistic induction signals—the “copying” and “pattern matching” operations (e.g., “copy the token that followed $A$ the last time $A$ appeared”). This information involves rapid transitions and occupies the high-frequency spectral band.
Because these signals occupy orthogonal frequency bands, they can coexist in the same “wire” (weight matrix, neuron) without destructive interference. This is precisely the principle underlying Spread Spectrum (CDMA) technology in telecommunications, where multiple signals share a single physical channel by occupying non-overlapping frequency or code spaces. Momentum Attention explicitly orthogonalizes these bands via the Symplectic-Filter Duality (Section 2.9), reducing the “interference” that manifests as polysemanticity in the spatial domain.
Remark 4.3 (From Spatial to Spectral Interpretability).
The Superposition phenomenon, viewed through the spectral lens, suggests a refinement of the interpretability program: rather than seeking to decompose the network into spatially localized “features per neuron,” we may achieve cleaner decomposition by analyzing features in the frequency domain. A neuron that appears “polysemantic” in the spatial domain may be cleanly “monosemantic” when its activations are decomposed into DC and AC spectral components. The Bode plot methodology (Algorithm 2) provides the practical tool for performing this spectral decomposition on trained attention heads.
4.3 The Broader Vision: From Statics to Dynamics
We wish to emphasize that our proposal is not a replacement for the MI program’s foundational circuit analysis, but rather a complementary extension. The computational graph discovered by Elhage et al. (2021) and Olsson et al. (2022) provides the topology of the circuit—which components connect to which, and what functions they compute. Our contribution adds the physics to this topology: conservation laws that govern how information flows through the circuit, and spectral dynamics that characterize how the circuit processes signals at different temporal scales.
The analogy to electrical engineering is instructive. An electrical circuit diagram (the computational graph) tells us the topology: which resistors, capacitors, and inductors are connected. But understanding the circuit’s behavior requires Kirchhoff’s Laws (conservation of charge and energy) and frequency-domain analysis (Bode plots, transfer functions). The MI program has given us the Transformer’s circuit diagram. We humbly propose that Hamiltonian mechanics and signal processing provide the Kirchhoff’s Laws and Bode analysis needed to understand the circuit’s dynamics.
This perspective reframes several outstanding puzzles:
The Hydra Effect is not a failure of the circuit abstraction, but Kirchhoff’s Current Law operating in the residual stream: total spectral current is conserved, so removing one path redistributes flow through others.
Superposition is not irreducible spatial interference, but frequency-division multiplexing: DC and AC signals share a channel without interference because they occupy orthogonal spectral bands.
The $L\geq 2$ depth constraint for induction (Olsson et al., 2022) is not a fundamental computational limit, but a consequence of the standard architecture’s “DC-coupled” design: depth serves as a proxy for derivative computation, which can be provided directly via momentum augmentation.
We believe this dynamic interpretability paradigm—analyzing the “resonance” rather than just finding the “circuit”—offers a productive path forward. A low-pass filter is a low-pass filter, regardless of which specific neurons implement it. By adopting the tools of Hamiltonian mechanics and signal processing, the interpretability community gains a robust, mathematically rigorous language for describing emergent phenomena that resist purely static decomposition.
5 Related Work
Mechanistic Interpretability. Our work builds directly on the foundational circuit analysis of Elhage et al. (2021) and Olsson et al. (2022), who established the rigorous framework for understanding transformers as computational graphs. We extend the geometric analysis of induction heads recently proposed by Musaf and others (2025) and the associative recall analysis by Sanford and Hsu (2024). Our spectral tools complement the automated circuit discovery methods of Conmy et al. (2023) and Olah et al. (2020), as well as the dictionary learning approaches of Bricken and others (2023) and the polysemanticity analysis of Goh and others (2021). The “circuit” metaphor has proven remarkably productive; our contribution extends this metaphor from static computational graphs to dynamic physical circuits.
Self-Repair and Dynamic Phenomena. The discovery of the “Hydra Effect” by McGrath et al. (2023)—where ablating one attention head causes other heads to spontaneously compensate—revealed that transformer computations exhibit a form of emergent self-repair that challenges purely static circuit decompositions. The finding that language model layers are “loosely coupled,” with ablations to one layer affecting only a small number of downstream layers, aligns naturally with our conservation-law framework: the symplectic structure predicts that spectral functions are distributed properties of layers rather than localized properties of individual heads. Similarly, the “Superposition” phenomenon identified by Elhage et al. (2022)—where neurons encode multiple features simultaneously—finds a natural resolution in our frequency-domain framework, where DC (semantic) and AC (mechanistic) signals can coexist in the same weight matrix without interference by occupying orthogonal spectral bands. We view our work as a complementary analytical toolkit that bridges these important empirical observations with the mathematical machinery of Hamiltonian dynamics and signal processing.
Frequency-Domain Approaches to Attention. Recent work has recognized the low-pass filtering behavior of self-attention and proposed various remedies. Most notably, Chen et al. (2025) introduce Frequency-Dynamic Attention Modulation (FDAM), which employs Attention Inversion to derive a complementary high-pass filter by algebraically inverting the attention matrix ($H_{\text{high}}\approx(I-A)x$). While FDAM achieves impressive results on dense prediction tasks in Vision Transformers, our approach differs fundamentally in its theoretical grounding. FDAM’s inversion is phenomenological—an empirically motivated “patch” applied to a leaky system. In contrast, our Momentum Attention derives the high-pass complement from first principles: the symplectic structure provides a conservation law (Liouville’s Theorem) that forbids the phase space collapse responsible for high-frequency signal loss in the first place. Where FDAM unblocks a clogged pipe, our framework constructs a pipe that cannot clog. This distinction has practical consequences: the symplectic constraint guarantees spectral energy is redistributed rather than created or destroyed, avoiding the potential for destructive interference or noise amplification that unconstrained high-pass augmentation might introduce.
Physics-Inspired Machine Learning. The intersection of Hamiltonian mechanics and deep learning has been explored extensively in Hamiltonian Neural Networks (Greydanus et al., 2019; Toth et al., 2020) and nonequilibrium thermodynamics (Sohl-Dickstein et al., 2015). We specifically draw inspiration from the renormalization group mappings by Mehta and Schwab (2014) and the statistical mechanics frameworks of Bahri et al. (2020) and Bondesan and Welling (2019). Lagrangian approaches by Cranmer and others (2020) and constrained optimization by Finzi and others (2020) also offer valuable perspectives on conservation laws in learning. Our work differs in applying these principles to the attention mechanism itself rather than to the overall network dynamics.
Signal Processing in Transformers. The role of positional encodings as filters has been studied in RoPE (Su and others, 2024) and ALiBi (Press et al., 2021), as well as Transformer-XL (Dai and others, 2019). Recent work by Kazemnejad and others (2024) and the “KV-Shifting” hypothesis by Hooper and others (2024) align with our kinematic findings. Our spectral forensics approach formalizes these observations using classical signal processing tools (Oppenheim and Willsky, 1996; Proakis and Manolakis, 2001; Kalman, 1960; Feynman et al., 1963). Efficient attention mechanisms like Reformer (Kitaev et al., 2020) and H2O (Zhang and others, 2024) also implicitly touch upon spectral sparsity. The Bode plot methodology we introduce provides a principled way to analyze any attention mechanism’s frequency response.
6 Conclusion
In this work, we have explored the potential of enriching the Transformer’s computational graph with physical conservation laws. By introducing Momentum Attention, we have shown that a simple symplectic augmentation ($p_{t}=q_{t}-q_{t-1}$) can imbue the model with fundamental conservation laws and spectral properties.
This intervention bridges the gap between Hamiltonian mechanics and signal processing. We have demonstrated that the “Symplectic Shear” is mathematically dual to a “High-Pass Filter,” unlocking powerful new tools for analysis—most notably Spectral Forensics. This framework not only explains why the model works (via the Orthogonality Theorem and Escape Routes) but also how to improve it. We humbly offer this work as an invitation for the community to apply the full arsenal of control theory and signal processing to the challenge of interpretability, extending the powerful “circuit” metaphor into the domain of physical dynamics.
Limitations and Future Work. While our experiments validate the theoretical predictions across 5,100+ controlled runs documented in Appendices A–R (with Appendix Q providing complete experimental configuration matrices), several limitations warrant discussion. First, the scale of our models (4M–350M parameters) remains modest compared to frontier systems; extending this framework to billion-parameter scales remains important future work, though the theoretical foundations are scale-agnostic. Second, the optimal momentum coupling $\gamma$ may vary across tasks and architectures—we provide extensive sweeps but acknowledge that adaptive coupling strategies warrant investigation; the attenuated scaling law $\gamma^{*}=4.17\times N^{-0.74}$ discovered in the Addendum to Appendix D provides initial guidance for this. Third, our spectral forensics methodology assumes access to model internals; applying these techniques to black-box models would require additional probing methods.
The symplectic structure naturally suggests extensions to other modalities (vision, audio, video) where temporal dynamics are more explicitly encoded. For video understanding, the kinematic prior could capture motion directly rather than requiring the model to infer it from static frames. For audio, the high-pass filtering interpretation connects to well-understood signal processing principles for speech recognition.
Reproducibility Statement. All experiments are fully reproducible via the 27 Jupyter notebooks provided in the supplementary material (Appendices A–R, including the Addenda to Appendices B, D, and E), with the complete notebook collection also available in the accompanying CODE-NOTEBOOKS-ARXIV.zip archive. Each notebook contains pre-embedded outputs from 5,169+ total experiments, enabling verification without GPU re-execution. Hardware configurations, hyperparameters, random seeds, and training details are documented in exhaustive detail following the “epistemic chronology” philosophy—preserving even productive failures and hypothesis revisions for complete scientific transparency.